# On the number of linear uniform hypergraphs with girth constraint

### v1

@@ 4 Analysis of the high-girth process
$\mathbbm{1}_{\{k\geqslant 1\}}\cdot|\Lambda_{\boldsymbol{f},L,k-1}(i)|$ $=O\bigl(n^{\alpha^{\prime}+(r-1)(e_{L}-k)-2}(nt_{{M}})^{k-1}\bigr)$ (4.15)
hold for any $\boldsymbol{f}\in{Q}(i)$, $L\in\mathcal{L}$ and $L^{\prime}\in{W}_{\boldsymbol{f},L,k}(i)$ with $0\leqslant k\leqslant e_{L}-2$, where $\alpha^{\prime}\in(\alpha_{0},\alpha)$. On the condition of the event $\mathcal{G}(i)$, the event $\mathcal{A}(i)$ holds with probability $1-o(n^{-\mu})$, that is,
@@
Since $\tilde{\xi}_{i}=O\bigl((nt)^{\ell-2}\bigr)$ in (3.12), q{q} in (3.19), ${y}_{m}$ in (3.20), ${w}_{L,k}$ in (3.21) and $\hat{\epsilon}_{y_{m}}$ in (3.28), we have
@@
$-\bigl(w_{L,k}+O(\epsilon_{w_{L,k}})\bigr)\biggl[\bigl(e_{L}-k-1\bigr)\biggl(\frac{r^{2}(r-1)^{2}}{2n^{2}p_{i}}+\frac{\tilde{\xi}_{i}}{n^{2}}\biggr)+O\Bigl(\sigma_{i}^{\ell}n^{\alpha-3}(nt_{{M}})^{\ell-2}\Bigr)\biggr]$
$=$ $-w_{L,k}\bigl(e_{L}-k-1\bigr)\biggl(\frac{r^{2}(r-1)^{2}}{2n^{2}p_{i}}+\frac{\tilde{\xi}_{i}}{n^{2}}\biggr)+O\Bigl(\sigma_{i}^{\ell}n^{\alpha+(r-1)(e_{L}-k)-r-3}(nt_{{M}})^{k+\ell-2}\Bigr).$ (4.17)
@@
The second term on the right side of the equation in (4.16) is reduced to
$\biggl[\bigl(e_{L}-k\bigr)\frac{{w}_{L,k-1}}{{q}}\pm\frac{2r!(e_{L}-k)}{p_{i}^{\binom{r}{2}}\xi_{i}}\beta_{k-1}\sigma_{i}^{k+1}n^{\alpha+(r-1)(e_{L}-k)-r-2}(nt_{\textsc{M}})^{k-1}\biggr]$
$=$ $\biggl[\bigl(e_{L}-k\bigr)\frac{{w}_{L,k-1}}{{q}}\pm\beta_{k}\sigma_{i}^{k+1}n^{\alpha+(r-1)(e_{L}-k)-r-2}(nt_{\textsc{M}})^{k-1}\biggr].$ (4.18)
@@
with $t_{{M}}$ in (3.26) and $\lambda>\frac{\ell}{\ell-1}$ in (3.27). For $k=0$, we only have the error term in (4.17), which is $O(\sigma_{i}^{\ell}n^{\alpha+(r-1)e_{L}-r-3}(nt_{{M}})^{\ell-2})$. ∎
@@
Next consider $|\Delta{Y}_{\boldsymbol{f}_{m}}(i)|$ for any $\boldsymbol{f}_{m}\in{K}_{m}(i)$ with $2\leqslant m\leqslant r-1$, and $|\Delta{W}_{\boldsymbol{f},L,k}(i)|$ for any $\boldsymbol{f}\in{Q}(i)$, $L\in\mathcal{L}$ and $0\leqslant k\leqslant e_{L}-2$. Let $\boldsymbol{e}_{i+1}\in{Q}(i)$ denote the available $r$-set that added to $\mathbb{H}(i)$. In fact, based on the analysis in Section 4.1.1 and Section 4.1.2, we have the ways to bound the one-step changes of $\Delta{Y}_{\boldsymbol{f}_{m}}(i)$ and $\Delta{W}_{\boldsymbol{f},L,k}(i)$. We formally characterize these ways into three boundedness parameters in higher uniformities below.
Define $\Pi_{\boldsymbol{f}_{m},\boldsymbol{e}_{i+1}}(i)$ as the collection of available $r$-sets in ${Y}_{\boldsymbol{f}_{m}}(i)$ that become unavailable due to the addition of the available edge $\boldsymbol{e}_{i+1}$ into $\mathbb{H}(i)$, where $|\boldsymbol{e}_{i+1}\cap\boldsymbol{f}_{m}|\leqslant 1$ because $\boldsymbol{f}_{m}\in K_{m}(i+1)$. Thus, we have
@@
Assuming the event $\mathcal{G}(i)\cap\mathcal{B}(i)$ holds, we now have $|\Delta{Y}_{\boldsymbol{f}_{m}}(i)|$ and $|\Delta{W}_{\boldsymbol{f},L,k}(i)|$ as follows.
@@
$\Delta{Y}^{\pm}_{{\boldsymbol{f}_{m}}}(i)$ $=\pm\biggl[\Delta{Y}_{{\boldsymbol{f}_{m}}}(i)-\frac{{y}_{m}^{\prime}(t)}{n^{2}}\biggr]-\frac{\epsilon_{y_{m}}^{\prime}(t)}{n^{2}}+O\biggl(\sup_{\varsigma\in[t,t+\frac{1}{n^{2}}]}\frac{|{y_{m}}^{\prime\prime}(\varsigma)|+|\epsilon_{y_{m}}^{\prime\prime}(\varsigma)|}{n^{4}}\biggr).$ (4.24)
@@
We have obtained that $\epsilon_{y_{m}}^{\prime}=\frac{n^{\alpha+r-m+1}}{n^{\alpha}+n^{2}t}$ by (3.23), and it is checked that the equation in (4.25) is true, where $t_{{M}}$ in (3.26) and $\lambda>\frac{\ell}{\ell-1}$ in (3.27).
@@
Claim 4.2.2 Assuming the event $\mathcal{G}(i)\cap\mathcal{B}(i)$ holds,
@@
Following the equation in (4.24), we indeed have shown
@@
$\Delta{W}^{\pm}_{\boldsymbol{f},L,k}(i)$ $=\pm\biggl[\Delta{W}_{\boldsymbol{f},L,k}(i)-\frac{{w}_{L,k}^{\prime}(t)}{n^{2}}\biggr]-\frac{\epsilon_{w_{L,k}}^{\prime}(t)}{n^{2}}+O\biggl(\sup_{\varsigma\in[t,t+\frac{1}{n^{2}}]}\frac{|{w}_{L,k}^{\prime\prime}(\varsigma)|+|\epsilon_{w_{L,k}}^{\prime\prime}(\varsigma)|}{n^{4}}\biggr).$ (4.26)
@@
Differentiating $\epsilon_{w_{L,k}}$ in (3.24) with respect to $t$, we have
$\epsilon_{w_{L,k}}^{\prime}$ $=(k+2)\beta_{k}\sigma_{i}^{k+1}n^{\alpha+(r-1)(e_{L}-k)-r+1}(n^{\alpha}+n^{2}t)^{-1}(nt_{{M}})^{k},$
and it is checked that the equations shown in (4.27) and (4.28) are true.
Finally, it is natural to obtain $\mathbb{E}[|\Delta{W}_{\boldsymbol{f},L,k}^{\pm}(i)|\,|\mathcal{F}_{i}]=O\bigl(\frac{|\epsilon_{w_{L,k}}^{\prime}|}{n^{2}}\bigr)$. ∎
Claim 4.2.4 Assuming the event $\mathcal{G}(i)\cap\mathcal{B}(i)$ holds,
@@
Following the equation shown in (4.26), we indeed have shown
@@
Let the stopping time $\tau_{\boldsymbol{f}_{m}}$ be the minimum of M{M}, the last step $i\geqslant 1$ such that $\boldsymbol{f}_{m}\in{K}_{m}(i)$, and the minimum $i\geqslant 1$ that the event $\bar{\mathcal{G}}^{+}({i})$ holds. Similarly, let the stopping time $\tau_{\boldsymbol{f}}$ be the minimum of M{M}, the last step $i\geqslant 1$ such that $\boldsymbol{f}\in{Q}(i)$, and the minimum $i\geqslant 1$ that the event $\bar{\mathcal{G}}^{+}({i})$ holds. It follows that
@@
In order to bound the above probability, the following tail probability formula in [12, 16, 7, 24] is applied to each inequality in (4.33). The following tail probability is due to Freedman [12], which is called to be Freedman’s martingale concentration inequality and was originally stated for martingales, while the proof for supermartingales is the same. It is also applied to study the existence of sparse Steiner triple systems to approximately answer a conjecture of Erdős in the previous works of Glock, Kühn, Lo and Osthus [16] and of Bohman and Warnke [7]. The simplified version of the Freedman’s martingale concentration inequality here [7, 16, 24] improves the Azuma-Hoeffding inequality when the expected one-step changes are much smaller than the worst case ones, that is, $\mathbb{E}[|\Delta{S}(i)|\,|\mathcal{F}_{i}]\ll\max_{i\geqslant 0}|\Delta{S}(i)|$.
@@
$\mathbb{P}\bigl[{Y}_{\boldsymbol{f}_{m}}^{\flat}(i\wedge\tau_{\boldsymbol{f}_{m}})\geqslant 0\text{ for some }i\geqslant 1\bigr]=n^{-\omega(1)}.$ (4.34)
@@
$\mathbb{P}\Bigl[{W}_{\boldsymbol{f},L,k}^{\flat}(i\wedge\tau_{\textbf{f}})\geqslant 0\text{ for some }i\geqslant 0\Bigr]=n^{-\omega(1)}.$ (4.35)
@@
In this section, for an integer $i$ with $0\leqslant i<{M}$, we still assume that the event $\mathcal{G}(i)$ holds, and prove Claim 4.1.3 and Claim 4.1.6.

### v2

@@ 4 Analysis of the high linear-girth process
$\mathbbm{1}_{\{k\geqslant 1\}}\cdot|\Lambda_{\boldsymbol{f},L,k-1}(i)|$ $=O\bigl(n^{\alpha^{\prime}+(r-1)(e_{L}-k)-2}(nt_{{M}})^{k-1}\bigr),$ (4.15)
hold for any $\boldsymbol{f}\in{Q}(i)$, $L\in\mathcal{L}$ and $L^{\prime}\in{W}_{\boldsymbol{f},L,k}(i)$ with $0\leqslant k\leqslant e_{L}-2$, where $\alpha_{0}=\frac{\ell-2}{\ell-1}$ and $\alpha^{\prime}\in(\alpha_{0},\alpha)$. On the condition of the event $\mathcal{G}(i)$, the event $\mathcal{A}(i)$ holds with probability $1-o(n^{-\mu})$, that is,
@@
Since $\tilde{\xi}_{i}=O((nt_{M})^{\ell-2})$ in (3.12), q{q} in (3.19), ${y}_{m}$ in (3.20), ${w}_{L,k}$ in (3.21) and $\hat{\epsilon}_{y_{m}}$ in (3.28), we have
@@
$-\bigl(w_{L,k}+O(\epsilon_{w_{L,k}})\bigr)\biggl[\bigl(e_{L}-k-1\bigr)\biggl(\frac{r^{2}(r-1)^{2}}{2n^{2}p_{i}}+\frac{\tilde{\xi}_{i}}{n^{2}}\biggr)\bigl(1-O(\hat{\epsilon}_{q})\bigr)+O\Bigl(\sigma_{i}^{\ell}n^{\alpha-3}(nt_{{M}})^{\ell-2}\Bigr)\biggr]$
$=$ $-w_{L,k}\bigl(e_{L}-k-1\bigr)\biggl(\frac{r^{2}(r-1)^{2}}{2n^{2}p_{i}}+\frac{\tilde{\xi}_{i}}{n^{2}}\biggr)+O\Bigl(\sigma_{i}^{\ell}n^{\alpha+(r-1)(e_{L}-k)-r-3}(nt_{{M}})^{k+\ell-2}\Bigr),$ (4.17)
where the term $\hat{\epsilon}_{q}\cdot(\frac{r^{2}(r-1)^{2}}{2n^{2}p_{i}}+\frac{\tilde{\xi}_{i}}{n^{2}})$ is absorbed into $O(\sigma_{i}^{\ell}n^{\alpha-3}(nt_{{M}})^{\ell-2})$ by $\tilde{\xi}_{i}=O((nt_{M})^{\ell-2})$ in (3.12) and $\hat{\epsilon}_{q}$ in (3.29), and then the term $\epsilon_{w_{L,k}}\cdot(\frac{r^{2}(r-1)^{2}}{2n^{2}p_{i}}+\frac{\tilde{\xi}_{i}}{n^{2}})$ is absorbed into
$O\bigl(\sigma_{i}^{\ell}n^{\alpha+(r-1)(e_{L}-k)-r-3}(nt_{{M}})^{k+\ell-2}\bigr)$
by $\tilde{\xi}_{i}$ in (3.12) and $\epsilon_{w_{L,k}}$ in (3.24). Specially,
$\epsilon_{w_{L,\ell-2}}\cdot\Bigl(\frac{r^{2}(r-1)^{2}}{2n^{2}p_{i}}+\frac{\tilde{\xi}_{i}}{n^{2}}\Bigr)=\Theta\bigl(\sigma_{i}^{\ell}n^{\alpha+(r-1)(e_{L}-k)-r-3}(nt_{{M}})^{k+\ell-2}\bigr)$
based on $\epsilon_{w_{L,k}}$ in (3.24).
@@
Hence,
$\frac{(e_{L}-k)\cdot{w}_{L,k-1}\cdot\hat{\epsilon}_{q}}{q}+\frac{2(e_{L}-k)\cdot\epsilon_{w_{L,k-1}}}{q}<\beta_{k}\sigma_{i}^{k+1}n^{\alpha+(r-1)(e_{L}-k)-r-2}(nt_{\textsc{M}})^{k-1}.$
It follows that the second term on the right side of the equation in (4.16) is reduced to
$\bigl(e_{L}-k\bigr)\frac{{w}_{L,k-1}}{{q}}\pm\beta_{k}\sigma_{i}^{k+1}n^{\alpha+(r-1)(e_{L}-k)-r-2}(nt_{\textsc{M}})^{k-1}.$ (4.18)
@@
is reduced to $\beta_{k}\sigma_{i}^{k+1}\gg\sigma_{i}^{\ell}(\log n)^{-\lambda(\ell-1)}$ by $t_{{M}}$ in (3.26) that is clearly true as $\sigma_{i}=\Theta(\log n)$ in (3.25) and $\lambda>\frac{\ell}{\ell-1}$ in (3.27).
For $k=0$, we have the only error term in (4.17), which is $O(\sigma_{i}^{\ell}n^{\alpha+(r-1)e_{L}-r-3}(nt_{{M}})^{\ell-2})$. ∎
@@
Next consider $|\Delta{Y}_{\boldsymbol{f}_{m}}(i)|$ for any $\boldsymbol{f}_{m}\in{K}_{m}(i)$ with $2\leqslant m\leqslant r-1$, and $|\Delta{W}_{\boldsymbol{f},L,k}(i)|$ for any $\boldsymbol{f}\in{Q}(i)$, $L\in\mathcal{L}$ and $0\leqslant k\leqslant e_{L}-2$. Let $\boldsymbol{e}_{i+1}\in{Q}(i)$ denote the available $r$-set that is added to $\mathbb{H}(i)$. In fact, based on the analysis in Section 4.1.1 and Section 4.1.2, we have the ways to bound the one-step changes of $\Delta{Y}_{\boldsymbol{f}_{m}}(i)$ and $\Delta{W}_{\boldsymbol{f},L,k}(i)$. We formally characterize these ways into three boundedness parameters in higher uniformities below.
Define $\Pi_{\boldsymbol{f}_{m},\boldsymbol{e}_{i+1}}(i)$ as the collection of (rm)(r-m)-sets in ${Y}_{\boldsymbol{f}_{m}}(i)$ that are removed due to the addition of the available edge $\boldsymbol{e}_{i+1}$ into $\mathbb{H}(i)$, where $|\boldsymbol{e}_{i+1}\cap\boldsymbol{f}_{m}|\leqslant 1$ because $\boldsymbol{f}_{m}\in K_{m}(i+1)$. Thus, we have
@@
Assuming the event $\mathcal{G}(i)\cap\mathcal{B}(i)$ holds, by Claim 4.1.6 and the equations shown in (4.19) and (4.20), we now have $|\Delta{Y}_{\boldsymbol{f}_{m}}(i)|$ and $|\Delta{W}_{\boldsymbol{f},L,k}(i)|$ as follows.
@@
$\Delta{Y}^{\pm}_{{\boldsymbol{f}_{m}}}(i)$ $=\pm\biggl[\Delta{Y}_{{\boldsymbol{f}_{m}}}(i)-\frac{{y}_{m}^{\prime}}{n^{2}}\biggr]-\frac{\epsilon_{y_{m}}^{\prime}}{n^{2}}+O\biggl(\sup_{\varsigma\in[t,t+\frac{1}{n^{2}}]}\frac{|{y_{m}}^{\prime\prime}(\varsigma)|+|\epsilon_{y_{m}}^{\prime\prime}(\varsigma)|}{n^{4}}\biggr).$ (4.24)
@@
As the equations of $\epsilon_{y_{m}}$ and $\sigma_{i}$ shown in (3.23) and (3.25), we have
$\epsilon_{y_{m}}^{\prime}=\frac{n^{\alpha+r-m+1}}{n^{\alpha}+n^{2}t}\ \text{and}\ \epsilon_{y_{m}}^{\prime\prime}=-\frac{n^{\alpha+r-m+3}}{(n^{\alpha}+n^{2}t)^{2}}.$
It is clearly correct that $\epsilon_{y_{m}}^{\prime}\gg\sup_{\zeta\in[t,t+\frac{1}{n^{2}}]}n^{-2}\cdot|\epsilon_{y_{m}}^{\prime\prime}(\zeta)|$. We further check
$\epsilon_{y_{m}}^{\prime}\gg\sigma_{i}^{\ell}n^{\alpha+r-m-1}(nt_{{M}})^{\ell-2}$
in (4.25) with $t_{{M}}$ in (3.26), which is true if $1\gg\sigma_{i}^{\ell}\cdot(\log n)^{-\lambda(\ell-1)}$ that holds as $\sigma_{i}=\Theta(\log n)$ in (3.25) and $\lambda>\frac{\ell}{\ell-1}$ in (3.27).
As the equation of ${y}_{m}$ shown in (3.20), differentiate ${y}_{m}$ with respect to $t$, and it is routine to check that $|{y}_{m}^{\prime}|=O(n^{r-m}(1+(nt)^{\ell-2}))=O(n^{r-m+\ell-2}t_{M}^{\ell-2})$ and $|{y}_{m}^{\prime\prime}|=O(n^{r-m+\ell-2}t_{M}^{\ell-3})$ such that
$\epsilon_{y_{m}}^{\prime}\gg\sup_{\zeta\in[t,t+\frac{1}{n^{2}}]}n^{-2}\cdot|{y}_{m}^{\prime\prime}(\zeta)|.$
@@
Claim 4.2.2 Assume the event $\mathcal{G}(i)\cap\mathcal{B}(i)$ holds. Then,
@@
Following the equation in (4.24) and the proof of Claim 4.2.1, we have shown
@@
$\Delta{W}^{\pm}_{\boldsymbol{f},L,k}(i)$ $=\pm\biggl[\Delta{W}_{\boldsymbol{f},L,k}(i)-\frac{{w}_{L,k}^{\prime}}{n^{2}}\biggr]-\frac{\epsilon_{w_{L,k}}^{\prime}}{n^{2}}+O\biggl(\sup_{\varsigma\in[t,t+\frac{1}{n^{2}}]}\frac{|{w}_{L,k}^{\prime\prime}(\varsigma)|+|\epsilon_{w_{L,k}}^{\prime\prime}(\varsigma)|}{n^{4}}\biggr).$ (4.26)
@@
Differentiating $\epsilon_{w_{L,k}}$ in (3.24) with respect to $t$, we have
$\epsilon_{w_{L,k}}^{\prime}$ $=(k+2)\beta_{k}\sigma_{i}^{k+1}n^{\alpha+(r-1)(e_{L}-k)-r+1}(n^{\alpha}+n^{2}t)^{-1}(nt_{{M}})^{k}.$
Now we check that the equations shown in (4.27) and (4.28) are true.
Firstly, we check
$\epsilon_{w_{L,0}}^{\prime}\gg\sigma_{i}^{\ell}n^{\alpha+(r-1)e_{L}-r-1}(nt_{{M}})^{\ell-2}$
in (4.27) for $k=0$, which is true if $2\beta_{0}\sigma_{i}\gg\sigma_{i}^{\ell}(\log n)^{-\lambda(\ell-1)}$ that holds as $\sigma_{i}=\Theta(\log n)$ in (3.25), $t_{M}$ in (3.26) and $\lambda>\frac{\ell}{\ell-1}$ in (3.27). For $k\geqslant 1$, the inequality in (4.27) is reduced to $(k+2)(n^{2}t_{{M}})>(n^{\alpha}+n^{2}t)$, which is clearly true as $\alpha\in(0,1)$ and $t\in[0,t_{M}]$ with $t_{M}$ in (3.26).
Taking the derivative of $\epsilon_{w_{L,k}}^{\prime}$ with respect to $t$ again, we have
$\epsilon_{w_{L,k}}^{\prime\prime}$ $=O\bigl(\sigma_{i}^{k+1}n^{\alpha+(r-1)(e_{L}-k)-r+3}(n^{\alpha}+n^{2}t)^{-2}(nt_{\textsc{M}})^{k}\bigr),$
which implies that $\epsilon_{w_{L,k}}^{\prime}\gg\sup_{\zeta\in[t,t+\frac{1}{n^{2}}]}n^{-2}\cdot{|\epsilon_{w_{L,k}}^{\prime\prime}(\zeta)|}$ in (4.28). Note that we can not change the power of $\sigma_{i}$ in the expression of $\epsilon_{w_{L,k}}$ defined in (3.24).
At last, we check $\epsilon_{w_{L,k}}^{\prime}\gg\sup_{\zeta\in[t,t+\frac{1}{n^{2}}]}n^{-2}\cdot{|{w}_{L,k}^{\prime\prime}(\zeta)|}$. Differentiating ${w}_{L,k}$ in (3.21) with respect to $t$, as the equations of $p_{i}$, $\xi_{i}$ and $\tilde{\xi}_{i}$ in (3.5), (3.9) and (3.12), it is checked that
${w}_{L,k}^{\prime}$ $=\mathbbm{1}_{\{k\geqslant 1\}}\cdot O\bigl(n^{(r-1)(e_{L}-k)+k-r}t^{k-1}\bigr)+O\bigl(n^{(r-1)(e_{L}-k)+k-r}(1+(nt)^{\ell-2})t^{k}\bigr).$
@@
${w}_{L,k}^{\prime\prime}$ $=O\bigl(n^{(r-1)(e_{L}-k)+\ell+k-r-2}t^{\ell+k-3}\bigr)$
$\quad+\mathbbm{1}_{\{k\geqslant 1\}}\cdot O\bigl(n^{(r-1)(e_{L}-k)+k-r}\bigl(1+(nt)^{\ell-2}\bigr)t^{k-1}\bigr)$
$\quad+\mathbbm{1}_{\{k\geqslant 2\}}\cdot O\bigl(n^{(r-1)(e_{L}-k)+k-r}t^{k-2}\bigr).$
Note that all these terms $n^{-2}\cdot{w}_{L,k}^{\prime\prime}$ for $t\in[0,t_{M}]$ are negligible compared to $\epsilon_{w_{L,k}}^{\prime}$ to complete the proof the inequality in (4.28).
Furthermore, it is natural to obtain $\mathbb{E}[|\Delta{W}_{\boldsymbol{f},L,k}^{\pm}(i)|\,|\mathcal{F}_{i}]=O\bigl(\frac{|\epsilon_{w_{L,k}}^{\prime}|}{n^{2}}\bigr)$. ∎
Claim 4.2.4 Assume the event $\mathcal{G}(i)\cap\mathcal{B}(i)$ holds. Then,
@@
Following the equation shown in (4.26) and the proof of Claim 4.2.3, we have shown
@@
Let the stopping time $\tau_{\boldsymbol{f}_{m}}$ be the minimum of M{M}, the last step $i\geqslant 1$ such that $\boldsymbol{f}_{m}\in{K}_{m}(i)$, and the maximum $i\geqslant 1$ that the event $\bar{\mathcal{G}}^{+}({i-1})$ holds. Similarly, let the stopping time $\tau_{\boldsymbol{f}}$ be the minimum of M{M}, the last step $i\geqslant 1$ such that $\boldsymbol{f}\in{Q}(i)$, and the maximum $i\geqslant 1$ that the event $\bar{\mathcal{G}}^{+}({i-1})$ holds. It follows that
@@
In order to bound the above probability, the following tail probability formula in [15, 19, 8, 32] is applied to each inequality in (4.33). The following tail probability is due to Freedman [15], which is called to be Freedman’s martingale concentration inequality and was originally stated for martingales, while the proof for supermartingales is the same. It is also applied to study the existence of sparse Steiner triple systems to approximately answer a conjecture of Erdős in the previous works of Glock et al. [19] and of Bohman and Warnke [8]. The simplified version of the Freedman’s martingale concentration inequality here [8, 19, 32] improves the Azuma-Hoeffding inequality when the expected one-step change is much smaller than the worst case one, that is, $\mathbb{E}[|\Delta{S}(i)|\,|\mathcal{F}_{i}]\ll\max_{i\geqslant 0}|\Delta{S}(i)|$.
@@
$\mathbb{P}\bigl[{Y}_{\boldsymbol{f}_{m}}^{\flat}(i\wedge\tau_{\boldsymbol{f}_{m}})\geqslant 0\text{ for some }i\geqslant 1\bigr]$
$\leqslant\mathbb{P}\bigl[{Y}_{\boldsymbol{f}_{m}}^{\flat}(i\wedge\tau_{\boldsymbol{f}_{m}})\geqslant{Y}_{\boldsymbol{f}_{m}}^{\flat}(0)+z\bigr]$
$\leqslant\exp\biggl[-\Omega\biggl(\frac{\log^{2}n\cdot n^{2\alpha+2r-2m-2}}{n^{\alpha^{\prime}+r-m-1}\cdot n^{\alpha+r-m-1}}\biggr)\biggr]$
$=n^{-\omega(1)}.$ (4.34)
@@
$\mathbb{P}\Bigl[{W}_{\boldsymbol{f},L,k}^{\flat}(i\wedge\tau_{\textbf{f}})\geqslant 0\text{ for some }i\geqslant 0\Bigr]$
$\leqslant\mathbb{P}\bigl[{W}_{\boldsymbol{f},L,k}^{\flat}(i\wedge\tau_{\textbf{f}})\geqslant{W}_{\boldsymbol{f},L,k}^{\flat}(0)+z\bigr]$
$\leqslant\exp\biggl[-\Omega\biggl(\frac{(\log n)^{k+2}\cdot n^{2\alpha+2(r-1)(e_{L}-k)-2r-2}(nt_{{M}})^{2k}}{n^{\alpha^{\prime}+(r-1)(e_{L}-k)-r-1}(nt_{{M}})^{k}\cdot n^{\alpha+(r-1)(e_{L}-k)-r-1}(nt_{{M}})^{k}}\biggr)\biggr]$
$=n^{-\omega(1)}.$ (4.35)
@@
In this section, assume that the event $\mathcal{G}(i)$ holds for an integer $i$ with $0\leqslant i<{M}$, and we prove Claim 4.1.3 and Claim 4.1.6.